OpenStudy (anonymous):

Please help :D will give medal and fan :) 1.Find four fractions that are equivalent to the fraction 2/3 2.Find four fractions that are equivalent to the fraction 8/5

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

your name i know someone named daniel

OpenStudy (anonymous):

4/6=x find x

OpenStudy (anonymous):

Do u understand now

OpenStudy (anonymous):

\[2/3=4/6=6/9=8/12\]